Dear Richard dear all,

please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: E. Pinna, A. Puglisi, L. Fini, M. Sozzi, P. Stefanini, S. Esposito, C. Arcidiacono, A. Riccardi, M. Xompero, R. Briguglio
MG: R. Biasi

Total # of person: 11

W unit

1) TT malfunctioning (more illumination of two pupils opposite to the center). Checking the images of the TT calibration we found that at 8V the TT mirrror can no longer produce a circular pattern. This behavior has been double checked with data taken already in Arcetri. We checked that the effect disappear when the TT unit is driven with less then 6V. The operation of pupil on sky centering on the CCD was performed in Arcetri with 6V modulation. Finally the 6V modulation level gives 5 times the maximum modulation we use in the AO loop.

2) Pupil geometry. The pupil geometry has been measured in Arcetri at a given modulation (6V@100Hz). Using this TT mirror set-up the pupil geometry measured is the correct one or 36.0 pixel side.

3) CCD47 cleaning. CCD47 has been unmounted form the WFS board. The CCD window internal to the camera has been cleaned. Final re-check of the CCD images has to be done.

4) ADC control The program controlling the Atmospheric dispersion correct has been upgraded to include a look up table giving the prisms counter rotation angle as a function of the telescope elevation as measured in the lab.

AO control SW

1) /etc/hosts files configured on both adsecdx and wfsdx.
2) Commissioning related documentation updated.
3) new IP assigned to ADSEC. The unit is fully reachable by adsecdx pc.
4) Marco Xompero start config adsec sw.

LBT672a

1) 1 crate BCU faulty breaking daisy-chain fast link communication. Fixed.
2) coil current check done
3) low-level safety checks done
4) the adsec fluxmeter is not working (reading zero flux). To be fixed.
5) AdSec supervisor SW compiled and configuration files set
6) AdSec supervisor SW test (startup and running) in progress


Logistic

1) Due to the snow on the mountain most of the Italian AO group left the telescope at 2:00PM. A. Riccardi, M. Xompero, R. Briguglio, R. Biasi remained at the telescope.

2) L. Fini and M. Sozzi left to Tucson and will be flying back to Europe today.


Next update tomorrow.

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback